OpenStudy (anonymous):

Help needed in related rates of change in the following problem: The area of a circle is decreasing at the rate of 2cm^2s^-1. How fast is the radius decreasing when the area is 9*pi I've tried finding the derivative of A with respect to r (dA/dr) and inversing it to get dr/dA didn't work either . Taking dr/dA directly didn't work either.

OpenStudy (phi):

\[ A= \pi r^2 \\ \frac{dA}{dr} = 2\pi r\\ dA= 2 \pi r \ dr \] from the given info we know \[ A= \pi r^2 \\ 9 \pi = \pi r^2 \] so we can find r we also know dr = 2 cm so we can find dA

OpenStudy (phi):

strictly speaking the ratio of infinitesimals such as \( \frac{dA}{dr} \) cannot be split like an ordinary fraction. what we do is say, the approximate differential is \[ \frac{\Delta A}{\Delta r}\] where we have a ratio of "tiny" but measurable quantities. Now it is ok to multiply both sides by \( \Delta r \) and we have the approximation \[ \Delta A \approx 2 \pi r \Delta r \]